Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
6x + y = 46
6x – y = 38

Answer by mananth(16949)

6 x + 1 y = 46 .............1
6 x -1 y = 38 .............2
Eliminate y
Add the two equations
12 x = 84
/ 12
x = 7
plug value of x in (1)
6 x + 1 y = 46
42 + 1 y = 46
1 y = 46 -42
1 y = 4
y = 4